### v2

# A Liquid-Nitrogen-Cooled ${}^{40}\mathrm{Ca}^{+}$ Ion Optical Clock with a Systematic Uncertainty of $4.4\times 10^{-19}$

Abstract
We report a single-ion optical clock based on the $4\mathrm{S}_{1/2}\rightarrow 3\mathrm{D}_{5/2}$ transition of the ${}^{40}\mathrm{Ca}^{+}$ ion, operated in a liquid nitrogen cryogenic environment, achieving a total systematic uncertainty of $4.4\times 10^{-19}$. It is the lowest reported uncertainty among all optical clocks to date. We employ a refined temperature evaluation scheme to reduce the frequency uncertainty due to blackbody radiation (BBR), and the 3D sideband cooling has been implemented to minimize the second-order Doppler shift. We have precisely determined the average Zeeman coefficient of the ${}^{40}\mathrm{Ca}^{+}$ clock transition to be $14.345(40)~{}\mathrm{Hz}/\mathrm{mT}^{2}$, thereby significantly reducing the quadratic Zeeman shift uncertainty. Moreover, the cryogenic environment enables the lowest reported heating rate due to ambient electric field noise in trapped-ion optical clocks.
Introduction—Optical frequency is currently the physical quantity with the highest measurement precision. Over the past two decades, with the advancement of atomic and optical physics theories as well as laser technology, optical clocks have undergone substantial improvements. At present, many optical clocks have achieved uncertainties at the order of  $10^{-18}$ or even $10^{-19}$ [Aeppli2024, marshall2025, Brewer2019, Ushijima2015, Huang2022, Tofful2024, Dorscher2021, zhiqiang2023, Li2024], enabling optical clocks to play an irreplaceable role in the field of precision measurement, such as investigating whether the fine-structure constant $\alpha$ varies over time [Godun2014], testing Lorentz invariance [Lange2021, Sanner2019], and detecting gravitational waves and dark matter[Kolkowitz2016, Filzinger2023]. Meanwhile, the ultra-high precision of optical clocks is driving the redefinition of the SI second [Dimarcq_2024] and has further promoted their applications in navigation and geodesy [takamoto2020, McGrew2018, meh2018].
The $\mathrm{Ca}^{+}$ ion features a simple energy level structure and requires relatively fewer numbers of lasers with weaker power. Moreover, it has a magic trapping RF (radio frequency) frequency that one can cancel the micromotion-induced Doppler and Stark shifts, making it one of the ideal optical clock candidates approaching very low total systematic uncertaitny[Huang2019]. The main contributor to the systematic uncertainty of the ${}^{40}\mathrm{Ca}^{+}$ ion optical clock is the BBR shift. To address this, we previously utilized a liquid-nitrogen cooling scheme to reduce the BBR shift uncertainty to $2.7\times 10^{-18}$[Huang2022]. However, due to the accuracy of the temperature evaluation, the BBR shift remains the largest contributor to the total systematic uncertainty[Huang2022]. Additionally, the second-order Doppler shift arising from secular motion is also one of the main factors restricting the further reduction of the total systematic uncertainty of the ${}^{40}\mathrm{Ca}^{+}$ ion optical clock to the $10^{-19}$ level.
In this Letter, we report on the research progress of the second-generation liquid-nitrogen cryogenic ion optical clock $\mathrm{Ca}^{+}\!\text{-}2$ (LNCIOC $\mathrm{Ca}^{+}\!\text{-}2$), which has achieved a total systematic uncertainty of $4.4\times 10^{-19}$. This is mainly attributed to our improvements in the design and installation scheme of the BBR cavity and ion trap, the incorporation of 3D sideband cooling, and the precise measurement of the quadratic Zeeman coefficient for the ${}^{40}\mathrm{Ca}^{+}$ ion’s clock transition.
The setup is illustrated in Fig. S1(a), and details of the trap device can be found in the references [Chou2010, Huang2022, Zeng2023]. The trap RF frequency is set at 24.801 MHz, which is close to the magic trapping frequency. A DC voltage of 160 V is applied to both endcap electrodes. To separate the two secular motion frequencies in radial directions thus enabling effective sideband cooling for all 3 motional modes, we apply DC voltages of ±$\pm 8$ V to the two RF coupled electrodes. Additionally, RF amplitude stabilization is introduced to stabilize the radial secular motion frequencies, with a long-term drift of less than 2 kHz. The excess micromotion of the ion is evaluated through the sideband spectra observed by three mutually orthogonal clock laser beams (beam 1, 2, and 3 in Fig. S1(b)) and is suppressed by optimizing the voltages applied to the compensation electrodes. The vacuum system of the liquid-nitrogen cryogenic optical clock is surrounded by four layers of magnetic shielding to reduce the magnetic field noise[Zeng2023]. All lasers used in LNCIOC $\mathrm{Ca}^{+}\!\text{-}2$ are shown in Fig. S1(b) The quantization axis is along the $y$ axis direction with a magnetic field of approximately 37.2 $\mu$T.
The pulse sequence begins with a 1.5-ms-long Doppler cooling. Then the ${}^{40}\mathrm{Ca}^{+}$ ion is optically pumped into one of the two sub-levels of the $\mathrm{S}_{1/2}$ state using the 397 nm $\sigma^{+}/\sigma^{-}$ lasers, achieving a state-preparation efficiency exceeding 99% within a 5-$\mu$s pulse length. Subsequently, the ${}^{40}\mathrm{Ca}^{+}$ ion is further cooled to near the 3D motional ground state through resolved sideband cooling (10 ms). To avoid the AC Stark shift, all lasers except the clock laser are blocked using mechanical shutters during the clock interrogation. The clock transition is probed with a hyper Ramsey scheme, employing a $\pi$-pulse duration of 4 ms and a free evolution time of 80 ms. Finally, a 2-ms-long state detection pulse is performed to determine the ion’s final state.
During the clock operation, we alternately lock the clock laser onto three pairs of Zeeman transitions: $\left|\mathrm{S}_{1/2},m=\mp 1/2\right\rangle\rightarrow\left|\mathrm{D}_{5/2},m=\mp 3/2\right\rangle$, $\left|\mathrm{S}_{1/2},m=\pm 1/2\right\rangle\rightarrow\left|\mathrm{D}_{5/2},m=\mp 3/2\right\rangle$, and $\left|\mathrm{S}_{1/2},m=\mp 1/2\right\rangle\rightarrow\left|\mathrm{D}_{5/2},m=\mp 5/2\right\rangle$, to cancel out the first-order Zeeman shift, the electric quadrupole shift, and the tensor Stark shift [Dube2014].
Blackbody radiation shift—The BBR shift can be expressed as [Porsev2006]
$\Delta\nu_{\mathrm{BBR}}=-\frac{1}{2}\left(831.9~{}\mathrm{V\,m^{-1}}\right)^{2}\left(\frac{T(\mathrm{K})}{300}\right)^{4}\Delta\alpha_{0}\left(1+\eta\right),$
where $T$ is the ambient temperature, $\Delta\alpha_{0}$ is the differential static polarizability for the ${}^{40}\mathrm{Ca}^{+}$ ion’s clock transition, and $\eta$ represents a minor dynamic correction that arises from the energy-level structure of the ${}^{40}\mathrm{Ca}^{+}$ ion and the spectral distribution of the BBR field. Compared to room-temperature optical clocks, the liquid-nitrogen cryogenic environment reduces the ambient temperature by approximately a factor of 4, leading to a reduction of the BBR shift uncertainty by more than 200 times with a similar evaluated temperature uncertainty. Consequently, the uncertainty in the BBR shift caused by $\Delta\alpha_{0}$ and $\eta$ is on the order of $10^{-21}$.
Compared to the first-generation liquid-nitrogen cryogenic ${}^{40}\mathrm{Ca}^{+}$ ion optical clock, we have made several key improvements in the design and assembly of the BBR cavity, ion trap, and electrodes. As shown in Fig. S1, the ventilation holes in the BBR cavity—originally included to maintain vacuum consistency with the external chamber—are now sealed with oxygen-free copper covers to effectively block thermal radiation from the room-temperature environment. Additionally, to enhance temperature uniformity among the ion trap, ion trap holder, endcap electrodes, BBR cavity, and bottom of the Liquid-Nitrogen Cryogenic container(LNC), thin silver foils are inserted at the mechanical interfaces to enhance thermal contact and ensure their temperatures closely track that of the LNC. Furthermore, phosphor-bronze wires are employed as the conductors for both temperature sensors and electrodes. Before entering the BBR cavity, these wires are wound around the LNC several times to ensure their temperatures fully equilibrate with that of the liquid nitrogen. Collectively, these modifications significantly enhance the temperature uniformity of the ambient environment.
Figure 1: Temperature measurements of the liquid-nitrogen-cooled cryogenic ${}^{40}\mathrm{Ca}^{+}$ ion optical clock 2. The abrupt temperature drops observed in (a) are due to the periodic refilling of liquid nitrogen.
We used PT100 thermistors in a three-wire configuration to eliminate temperature measurement errors from wire resistance. All sensors were individually calibrated to an accuracy of 0.1 K over the 75–85 K temperature range. In LNCIOC $\mathrm{Ca}^{+}\!\text{-}2$, five sensors were employed to monitor the temperatures at the bottom of the LNC, the endcap electrodes, the ion trap holder, and the BBR cavity. See Fig. S1(a) in the supplementary material for the detailed locations of these components.
To directly assess the temperature of the ion trap, we constructed a replica system with an identical ion-trapping setup and BBR cavity to that of LNCIOC $\mathrm{Ca}^{+}\!\text{-}2$. An additional temperature sensor was attached to the ion trap within this replica system. The ion trap exhibits the highest temperature among all components, due to its direct connection to four electrodes. When the RF is switched on or off, a temperature change of approximately 0.4 K is observed. We took the upper and lower bounds of all 10 temperature sensors—spanning both the operational and replica systems—yielding a temperature estimate of $79.5\pm 1.4$ K. Considering the 0.1 K calibration uncertainty, the final BBR temperature is evaluated as $79.5\pm 1.5$ K, leading to a BBR uncertainty of $3.5\times 10^{-19}$, nearly an order of magnitude lower compared to the first-generation liquid-nitrogen cryogenic ${}^{40}\mathrm{Ca}^{+}$ optical clock.
Second-order Doppler shift—We implemented 3D sideband cooling in LNCIOC $\mathrm{Ca}^{+}\!\text{-}2$ to suppress the second-order Doppler shift. Due to small Lamb-Dicke parameters for the 3D secular motions, which are $\eta_{x}=0.025$, $\eta_{y}=0.027$, and $\eta_{z}=0.043$ respectively, the second-order sidebands are too weak to be experimentally resolved. Therefore, we sequentially selected three first-order red sidebands to reduce the occupation numbers. Detailed sequences are illustrated in Fig. S2.
Figure 2: Red- and blue-sideband Rabi oscillations for the three motional modes after sideband cooling. The orange dots represent the experimentally measured transition probabilities, and the blue solid lines are thermal-distribution fits to these data.
Figure 3: Measurement of the heating rates.Panels (a), (c), and (e) present red-sideband Rabi oscillations of the $x$-, $y$-, and $z$-motional modes, respectively, measured at different dark times following sideband cooling. By fitting these curves with thermal-distribution models, the mean occupation numbers at different delays are extracted. (b), (d), and (f) show linear fits to the data in (a), (c), and (e), respectively. The slope of each fit corresponds to the heating rate, and the intercept gives the mean occupation number after sideband cooling. All fit results include 95% confidence intervals as uncertainty estimates.
Table 1: Evaluation of the second-order Doppler shift for LNCIOC $\mathrm{Ca}^{+}\!\text{-}2$.
$x$ $y$ $z$
$\omega/2\pi$ (MHz) 3.531 3.166 2.580
$\eta$ 0.025 0.027 0.043
TD/quantum ($10^{-19}$) 3.7 3.3 2.4
$n_{0}$ (quanta) 0–0.03 0–0.04 0–0.11
$\dot{\bar{n}}$ (quanta/s) 0.43(7) 0.53(12) 1.31(43)
Total [$10^{-19}$] $-5.8\pm 0.4$
The result of 3D sideband cooling is shown in Fig. 2. Assuming that the Fock-state distribution is thermal after sideband cooling, we extract the mean occupation number, $\bar{n}$, by fitting the experimental data of the red-sideband Rabi oscillation to a thermal distribution model. Furthermore, as shown in Fig. 3, by inserting a ’dark time’ ranging from 30 to 140 ms between sideband cooling and red-sideband interrogation, we extracted the heating rates due to ambient electric field noise. This heating effect can rapidly drive the Fock state towards a thermal distribution, however, the Fock-state distribution after sideband cooling may not initially be thermal [Chen2017, Rasmusson2021]. To address the mean occupation number after sideband cooling, except using the assumption of thermal distribution, we also used the intercept obtained from a linear fit to the heating rate measurements.
Over a period of nearly one month, we carried out five repeated measurements of both the initial mean occupation numbers after sideband cooling and the heating rates. As presented in Fig. S4. Under an 80-ms interrogation time, the uncertainty in the second-order Doppler shift is evaluated to be $4\times 10^{-20}$. The specific data are shown in Table 1.
Second-order Zeeman shift—The output frequency of the ${}^{40}\mathrm{Ca}^{+}$ ion optical clock is determined by averaging over the three pairs of Zeeman transitions mentioned above. By comparing the output frequencies of two ${}^{40}\mathrm{Ca}^{+}$ ion optical clocks, we precisely measured the average second-order Zeeman coefficient $\alpha_{z}$ of the three pairs of Zeeman transitions. One clock served as a frequency reference, with its DC magnetic field fixed at a constant value $B_{0}$, and another served as the modulated clock with varing dc magnetic fields B. The frequency difference $\delta_{z}$ between the two clocks was measured as a function of $B$ and $B_{0}$, and can be described by the relation [Brewer2019]
$\delta_{z}=\alpha_{z}(B^{2}-B_{0}^{2}).$
By varying the DC magnetic field $B$ of the modulated clock, we obtained a set of frequency differences between the two clocks, as shown in Fig. S6. A linear fit to these data yielded $\alpha_{z}=14.345(40)~{}\mathrm{Hz/mT}^{2}$. The uncertainty of $\alpha_{z}$ was determined from the 95% confidence interval of the fit.
We derived the DC magnetic field of LNCIOC $\mathrm{Ca}^{+}\!\text{-}2$ from the measured Zeeman splittings, determining it to be $37.197(1)\ \mu\mathrm{T}$, corresponding to a second-order Zeeman shift uncertainty of $1.4\times 10^{-19}$. Although the DC magnetic field may exhibit significant drift over a long period, daily variations remain well within this uncertainty budget.
Furthermore, due to imperfections in the trap symmetry, the RF field generates an oscillating magnetic field at the ion’s position, with a frequency equal to that of the RF [Gan2018]. This RF-induced AC magnetic field can be characterized by measuring its effect on the ratio of Landé $g$-factors associated with the clock transition [Ma2024, Arnold2019]. The modulus squared of the $y$-component of the AC magnetic field generated by the RF along the DC magnetic field direction is evaluated to be
$\langle B_{y}^{2}\rangle=11(9)\times 10^{-14}~{}\mathrm{T}^{2}.$
This yields an estimated upper bound of $4\times 10^{-21}$ for the second-order Zeeman shift.
Excess micromotion (EMM) shift—There exists a magic drive frequency of $24.801(2)\ \mathrm{MHz}$ for the ${}^{40}\mathrm{Ca}^{+}$ ion clock transition, at which the EMM-induced second-order Doppler shift and the Stark shift cancel out [Huang2019]. The RF frequency was set to $24.801\ \mathrm{MHz}$ for LNCIOC $\mathrm{Ca}^{+}\!\text{-}2$ to minimize the EMM shift. Through long-term measurements of the sideband spectra in three mutually perpendicular directions, the frequency shift caused by EMM is evaluated to be $(0\pm 1.4)\times 10^{-19}$.
Background gas collisions—The collision-induced frequency shift (CFS) arising from interactions with background helium gas is evaluated using a nonperturbative quantum channel framework developed by A. C. Vutha  [Vutha2017]. This method has been validated in prior studies involving Sr+ [Vutha2017] and Al+  [Hankin2019, Davis2019] ions colliding with He and H2. As it is not feasible to measure the pressure inside the BBR cavity, we use the pressure measured in the ambient room-temperature environment outside the BBR cavity, which is $3\times 10^{-9}~{}\mathrm{Pa}$, as an over-estimate of the pressure inside the BBR cavity. The temperature of the background gas is $78.5~{}\mathrm{K}$. Subsequently, a particular calculation approach yields a collision frequency shift of $1.1\times 10^{-19}$. Details can be found in the Supplementary Material.
Other shifts—The hyper Ramsey excitation scheme was employed to suppress the AC Stark shift induced by the clock laser [Zhang2020] and the chirp effect of the acousto-optic modulator (AOM), ensuring that both frequency shifts are less than $1\times 10^{-19}$. The total systematic frequency shifts of LNCIOC $\mathrm{Ca}^{+}\!\text{-}2$ are presented in Table 2.
Table 2: Systematic shifts and uncertainties of the $\mathrm{LNC}\cdot{}^{40}\mathrm{Ca}^{+}$ion optical clock 2.
Effect Shift ($10^{-19}$) Uncertainty ($10^{-19}$)
BBR 45.5 3.5
Excess micromotion 0.0 1.4
Secular motion -5.8 0.4
Quadratic Zeeman (DC) 482.9 1.4
Quadratic Zeeman (AC) 0.0 0.04
Background gas 0.0 1.1
Residual 1st Zeeman 0.0 0.2
Electric quadrupole 0.0 0.1
AOM phase chirp 0.0 1.0
Clock laser Stark 0.0 0.6
Total 522.9 4.4
Conclusion—We have developed a second-generation liquid-nitrogen cryogenic ${}^{40}\mathrm{Ca}^{+}$ ion optical clock with an uncertainty of $4.4\times 10^{-19}$, which represents an improvement of approximately one order of magnitude compared to our previous result [Huang2022]. The cryogenic environment not only significantly reduces the blackbody radiation (BBR) shift and its uncertainty, but also markedly reduces the heating rates caused by ambient electric field noise. As the uncertainty and stability performance of ion optical clocks continue to improve, the heating effect has gradually become one of the main limiting factors for the system uncertainty and ion coherence time of optical clocks. Our work experimentally verifies the excellent suppression effect of the cryogenic environment on the heating rate, providing a feasible solution for higher-performance ion optical clocks.
For this liquid-nitrogen-cooled ${}^{40}\mathrm{Ca}^{+}$ ion optical clock, the BBR shift remains the primary contributor to the systematic uncertainty. Developing ${}^{40}\mathrm{Ca}^{+}$ ion optical clocks cooled with helium at even lower temperatures or employing more accurate temperature assessment schemes holds promise for further reducing the BBR shift to the order of $10^{-20}$. Meanwhile, by creating conditions to measure the ion exchange rate between two ions, one can precisely evaluate the vacuum level in directly cooled single-ion optical clocks. Additionally, developing a more comprehensive theoretical model for the collision frequency shift will enable further reduction of the background gas collision-induced frequency shift. In the next step, we will also enhance the performance of the clock laser and magnetic shielding to extend the clock interrogation time, aiming to achieve a ${}^{40}\mathrm{Ca}^{+}$ ion optical clock with a second-level stability at the $10^{-16}$ level. This, in turn, will allow the application of the $10^{-19}$-level precision of the ${}^{40}\mathrm{Ca}^{+}$ ion optical clock to precision measurements such as the frequency ratio between different optical clocks, the time variation of the fine-structure constant ($\alpha$), and the determination of millimeter-level height differences and tidal potentials.
I ACKNOWLEDGMENTS
We are deeply grateful to Dr. XXX for his valuable discussions, and to Professor XXX for their careful review of the manuscript. This work is supported by the National Key R&D Program of China (Grant No. 2022YFB3904001), Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0300901), the National Natural Science Foundation of China (Grant Nos. 12204494, 12121004, 12320101003 and 11934014), the CAS Youth Innovation Promotion Association (Grant No Y2022099), the Natural Science Foundation of Hubei Province (Grant Nos. 2022CFA013 and 2023EHA006), and the CAS Project for Young Scientists in Basic Research (Grant Nos. YSBR-055 and YSBR-085).